OpenStudy (anonymous):

A boy walked over hills at 3 kilometers per hour and on flat roads at 4 kilometers per hour. He walked 18 kilometers in 5 hours. How much of the trip was over hills?... I usually use some from of the formula d=rt or distance=rate*time Someone help!!! I already know that the answer is 6 kilometers, but don't know how to work it agebraically.

OpenStudy (anonymous):

3*2=6 (Km) 4*3=12 (Km) 12+6=18 (Km) 3+2=5 (Hrs)

OpenStudy (anonymous):

One question... where is the 3*2=6 coming from?